A new car is purchased for 20700 dollars. The value of the car depreciates at 11% per year. What will the value of the car be, to the nearest cent, after 11 years?


1
Expert's answer
2021-12-19T13:59:43-0500

 value after n years:

"D=P(1-r)^n"

where P = present value,

P = present value,

n = no. of years


"D=20700(1-0.11)^{11}=\$5744.60"